The average travel time from your residence to your school is 35 minutes with a standard deviation of 10 minutes. If you want to be 99% certain that you will not be late for first 9:00 am class, what is the latest time you should leave home? Assume that the time travel is normally distributed


1
Expert's answer
2021-12-21T02:35:43-0500
"P(Z<z)=0.99"

"z\approx2.3263"

"z=\dfrac{x-\mu}{\sigma}"

"\dfrac{x-35}{10}\approx2.3263"

"x=2.3263(10)+35"

"x=58.263 \ min"




"9\ hours-58.263\ min=8hours\ 1.737\ min"

The latest time you should leave home is "8:02."